# A New Probe of Cosmic Birefringence Using Galaxy Polarization and Shapes

### v1

@@ Abstract
where $s(z)=\int_{0}^{z}c/H(z^{\prime})\,{\rm d}z^{\prime}$ is the comoving distance to redshift $z$; $H(z)$ is the expansion rate for $H_{0}=70\,\mathrm{km\,s^{-1}\,Mpc^{-1}},\Omega_{\Lambda}=0.7,\Omega_{m}=0.3$; $\Omega$ is the surveyed sky area in steradians; $\phi$ is the $4.8\,\mathrm{GHz}$ radio luminosity function (per unit comoving volume) of spiral galaxies at $z=0$ (details can be found in App. D); $L_{4.8}^{\mathrm{max}}=10^{25}\,\mathrm{W\,Hz^{-1}}$ is a generous upper bound on the radio luminosity of a spiral galaxy; $L_{4.8}^{\mathrm{min}}$ is the luminosity of a galaxy at redshift $z$ with an observed spectral flux density of $S_{\nu}^{\mathrm{min}}$:
@@
and $\lambda(z)$ describes the luminosity evolution with $z$ for spiral galaxies [50]. The upper limit in the $z$ integral should not exceed $z=1.5$; above that, nearly all massive halos host spheroidal galaxies [57]. The lower limit in the $z$ integral ideally should exceed the redshift of detectable sources of birefringence. There is then a trade-off between the galaxy number and the accumulation of birefringence along the line of sight. A more sophisticated treatment would consider birefringence sources interspersed amongst the galaxies throughout all redshift $1.50<z<$ 1.5. We leave such model-dependent calculations to a future study.
The highest sensitivity achievable with available data in the near future will likely come from correlating polarized SKA radio galaxies with LSST shape data (or using radio shapes from resolved SKA images). Similar to Ref. [58], we consider Wide, Deep, and Ultradeep reference surveys, which have $\sigma_{I}=1\,\mathrm{\mu Jy}$, $0.2\,\mathrm{\mu Jy}$, and $0.05\,\mathrm{\mu Jy}$, and cover $1000\,\mathrm{deg}^{2}$, $20\,\mathrm{deg}^{2}$, and $1\,\mathrm{deg}^{2}$ on the sky, respectively, but for SKA-MID Band 5a (4.64.64.64.68.58.58.5\,8.5GHz [59]). Including galaxies whose integrated radio intensities have ${\rm SNR}>5$ at $4.8\,\mathrm{GHz}$, we estimate the number of measurable spiral galaxies at $1.50.2<z<$ 1.5 to be $N=1.6\times 10^{6}$, $2.1\times 10^{5}$ and $2.8\times 10^{4}$ for each survey, respectively. For $1.50.8<z<$ 1.5, the numbers are $N=2.8\times 10^{5}$, $9.5\times 10^{4}$ and $1.7\times 10^{4}$, respectively. App. D details the calculation behind these numbers.
While most detectable galaxies will have polarization measurement errors making an important contribution to the variance of the estimator $\hat{\alpha}$, they collectively provide a non-negligible amount of information to reduce the variance in the estimation of $\alpha$. Consider $N$ detectable spiral galaxies in a small sky patch ${\rm d}\Omega$, each with a rotation estimator $\hat{\alpha}_{i}$. From a general linear combination $\hat{\alpha}_{\mathrm{best}}\equiv\sum^{N}_{i=1}w_{i}\,\hat{\alpha}_{i}$, we can obtain the weights that lead to an unbiased minimum-variance estimator, namely $w_{i}=\expectationvalue{\hat{\alpha}_{i}^{2}}^{-1}/\sum_{j}\expectationvalue{\hat{\alpha}_{j}^{2}}^{-1}$. For different galaxies, $\expectationvalue{\hat{\alpha}_{i}^{2}}$ only depends on the unpolarized radio SNR, which in turn only depends on $z$ and $L_{4.8}$. Let us denote it as $\expectationvalue{\hat{\alpha}^{2}}(x)$, where $x=(z,\,\log L_{4.8})$.
With Eq. (18) rewritten as $N={\rm d}\Omega\,\int\Phi(x)\,{\rm d}x$, the weights take the form
@@
where $\Phi(x)$ quantifies the galaxy distribution with redshift and luminosity as defined in Eq. (40). This leads to
@@
Accounting for the SNR distribution is therefore equivalent to having $N$ identically measured galaxies, each with an effective estimator variance of
@@
While the $x$-integration formally depends on the choice of the threshold intensity $S_{\nu}^{\mathrm{min}}$ as in Eq. (18), the integral approaches a minimal value as $S_{\nu}^{\mathrm{min}}$ decreases, until it is practically saturated for $S_{\nu}^{\mathrm{min}}/\sigma_{I}\lesssim 5$.
In Fig. 3, the noise angular power spectrum of the rotation estimator is plotted, along with the corresponding sensitivity curves for current and projected CMB surveys [20]. SKA will enable a noise power spectrum lower than that of CMB-S4 for $L\gtrsim 5$, but probe a more limited $z$ range for the birefringence sources. The noise spectra are compared to the maximal signal allowed by Planck [20] due to a network of ultralight axion strings [13] approximated by the loop crossing model [60].
@@
Discussion — Using Eq. (11), noisy rotation estimations can be done at discrete galaxy locations $\{\hat{\alpha}_{i}\}_{i}$, which is suited for real-space approaches that exploit the non-Gaussian statistics of $\alpha(\bm{\mathbf{\hat{n}}})$. This is applicable to birefringence from an axion string network [61, 62].
Measuring polarization rotation in the CMB has the advantage of being sensitive to all birefringence sources up to $z\approx 1100$. Spiral galaxies do not effectively probe birefringence sources at $z\gtrsim 1.5$. Nonetheless, galaxies present an opportunity for extracting tomographic information, i.e. the $z$-dependence of the birefringence sources. One could compare rotation estimated from the CMB with that from radio galaxies, or compare estimates from galaxies in different $z$ bins. For example, if axion strings exist for much of the cosmic history, this would enable the measurement of their size and number evolution with $z$ [60]. Furthermore, an axion field that is homogeneous but dynamic for $z\lesssim 1100$ is indistinguishable from a constant uniform axion field via a CMB measurement, but could show up as $z$-dependent in galaxy-based birefringence tomography. This also means that any detection of cosmic birefringence through the CMB [14] can then be validated through galaxy surveys. Foreground systematics that are unaccounted for, e.g. from the MW, would show up coherently for galaxies in every redshift bin and could thus be identified and subtracted.
Conclusion — We have proposed a new statistical technique to estimate anisotropic cosmic birefringence from correlation between radio polarization and shape in spiral galaxies. This offers an independent technique that complements CMB probes. This will be enabled by a synergy between upcoming large radio galaxy surveys such as SKA and optical imaging surveys such as LSST. The noise floor of the polarization rotation angular power spectrum of the galaxy-based method using SKA will be better than that of CMB-S4. We plan to test the empirical correlations of [24] on cosmological galaxy samples using data from existing smaller radio surveys [35, 48, 42].
@@ Acknowledgements.
We thank Neal Dalal, Marc Kamionkowski, Dustin Lang, Jessie Muir and Kendrick Smith for useful discussions and comments. L.D. acknowledges research grant support from the Alfred P. Sloan Foundation (Award Number FG-2021-16495), and support of Frank and Karen Dabby STEM Fund in the Society of Hellman Fellows. S.F. is supported by Lawrence Berkeley National Laboratory and the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade.
@@ Appendix A Quadratic estimator
Polarization and shape measurements have errors. For example, the polarized flux has a squared error
$\sigma_{P}^{2}=\frac{\sigma_{I}^{2}}{2},$ (24)
where $\sigma_{I}$ is the RMS flux of the detector in units of Jansky. We assume that the effect of measurement errors on the variance of the quadratic estimator $\hat{\alpha}$ is dominated by the measurement error for polarization, and the shape error is much smaller. In this case,
$\tilde{\sigma}_{X,Y}^{2}=\sigma_{X,Y}^{2}+\frac{\sigma_{I}^{2}\,\sigma_{\epsilon}^{2}}{2\,I^{2}}=\sigma_{X,Y}^{2}+\frac{\sigma_{\epsilon}^{2}}{2\,\mathrm{SNR}^{2}},$ (25)
where the total flux SNR enters the expression.
@@
where the coefficients are to be determined. Under a polarization rotation $\alpha$ at the galaxy’s location, the observed quantities are transformed:
$\tilde{X}\to\tilde{X}-2\alpha\,Y,\quad\tilde{Y}\to\tilde{Y}+2\alpha\,X.$ (27)
Note that the transformation is proportional to quantities without measurement noises, since cosmic birefringence is applied before detection. By requiring that the estimator $\hat{\alpha}$ be unbiased, namely $\expectationvalue{\hat{\alpha}}=\alpha$, we have
@@
Using these, $u$ can be expressed in terms of $u^{\prime}$ and $v^{\prime}$, and $v$ can be expressed in terms of $w$. It remains to minimize the variance of the estimator, $\expectationvalue{\hat{\alpha}^{2}}$, under the null hypothesis, with respect to $u^{\prime}$, $v^{\prime}$, and $w$. After some algebra, we obtain the claimed result in Eq. (11):
@@ Appendix B Maximum likelihood approach
where $\mathbf{N}_{i}$ is the measurement error of $\mathbf{X}_{i}$; $\overline{\mathbf{X}}_{i}$ is the mean of $\mathbf{X}$ over an ensemble of galaxies; $C_{\mathbf{X}_{i}}$ is the covariance matrix of the intrinsic, unrotated quantities $\mathbf{X}_{i}$ (which in general may depend on the type or other properties of the $i$-th galaxy); $C_{\mathbf{N}_{i}}$ is the covariance matrix of the measurement noise given by the second term in Eq. (10); and $M(\alpha(\bm{\mathbf{\hat{n}}}_{i}))$ is the effect of polarization rotation on the intrinsic quantities. Eq. (31) assumes that each $\mathbf{X}_{i}$ is a random Gaussian variable, the same approximation we used in the main text, meaning $C_{\mathbf{X}_{i}}=\operatorname{diag}(\tilde{\sigma}_{X_{i}}^{2},\tilde{\sigma}_{Y_{i}}^{2})$. Our goal is then to obtain an estimator $\hat{\alpha}_{\mathrm{ML}}(\bm{\mathbf{\hat{n}}})$ which maximizes $\ln\mathcal{L}$ given a set of observations $\{\mathbf{X}_{i}\}$.
@@
$\ln\mathcal{L}[\alpha;\mathbf{X},\mathbf{N}=\mathbf{N}_{\mathrm{max}}]=-\frac{1}{2}(\mathbf{X}-\mathbf{M}^{-1}\overline{\mathbf{X}})^{-1}(C_{\mathbf{N}}+\mathbf{M}^{T}C_{\mathbf{X}}\mathbf{M})^{-1}(\mathbf{X}-\mathbf{M}^{-1}\overline{\mathbf{X}}).$ (34)
@@
We insert this expansion into Eq. (34) and take the derivative with respect to $\alpha$, keeping terms up to linear order:
$\frac{\partial\mathcal{L}}{\partial\alpha}=\frac{2}{\tilde{\sigma}_{X}^{2}\,\tilde{\sigma}_{Y}^{2}}\left[\tilde{\sigma}_{Y}^{2}\,\overline{X}\,\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}\right]-\frac{4\,\alpha}{\tilde{\sigma}_{X}^{4}\,\tilde{\sigma}_{Y}^{4}}\left[(\cdots)+(\cdots)\tilde{X}+(\cdots)\tilde{X}^{2}+(\cdots)\tilde{Y}^{2}\right],$ (36)
where ()($\cdots$)( ⋯ ) are combinations of statistical quantities whose exact forms are not enlightening. This results in the maximum-likelihood estimator
$\hat{\alpha}_{\mathrm{ML}}=\frac{\tilde{\sigma}_{X}^{2}\,\tilde{\sigma}_{Y}^{2}\left[\tilde{\sigma}_{Y}^{2}\,\overline{X}\,\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}\right]}{2\left[(\cdots)+(\cdots)\tilde{X}+(\cdots)\tilde{X}^{2}+(\cdots)\tilde{Y}^{2}\right]}.$ (37)
@@
$\hat{\alpha}_{\mathrm{ML}}=\frac{\tilde{\sigma}_{X}^{2}\,\tilde{\sigma}_{Y}^{2}\left[\tilde{\sigma}_{Y}^{2}\,\overline{X}\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}\right]}{2\left[(\cdots)+(\cdots)\overline{X}+(\cdots)(\tilde{\sigma}_{X}^{2}+\overline{X}^{2})+(\cdots)\tilde{\sigma}_{Y}^{2}\right]}.$ (38)
After some simplification, the quadratic estimator in Eq. (11) is recovered:
$\hat{\alpha}_{\mathrm{ML}}=\frac{\tilde{\sigma}_{Y}^{2}\,\overline{X}\,\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}}{2\left[(\sigma_{X}^{2}-\sigma_{Y}^{2})^{2}+\tilde{\sigma}_{X}^{2}\,\overline{X}^{2}\right]},$ (39)
which agrees with the quadratic estimator Eq. (11). When $\alpha=0$, the variance $\expectationvalue{\hat{\alpha}_{\mathrm{ML}}^{2}}$ is therefore also given by Eq. (12).

### v2

@@ Abstract
where $s(z)=\int_{0}^{z}c/H(z^{\prime})\,{\rm d}z^{\prime}$ is the comoving distance to redshift $z$; $H(z)$ is the expansion rate for $H_{0}=70\,\mathrm{km\,s^{-1}\,Mpc^{-1}},\Omega_{\Lambda}=0.7,\Omega_{m}=0.3$; $\Omega$ is the surveyed sky area in steradians; $\phi$ is the $4.8\,\mathrm{GHz}$ radio luminosity function (per unit comoving volume) of spirals at $z=0$ (see App. D); $L_{4.8}^{\mathrm{max}}=10^{25}\,\mathrm{W\,Hz^{-1}}$ is a generous upper bound on the radio luminosity of a spiral galaxy; $L_{4.8}^{\mathrm{min}}$ is the galaxy luminosity at redshift $z$ with an observed spectral flux density $S_{\nu}^{\mathrm{min}}$:
@@
for the radio spectral index $\alpha=-0.7$ [51, 52], and $\lambda(z)$ describes the luminosity evolution with $z$ for spiral galaxies [50]. The upper limit in the $z$ integral should not exceed $z=1.5$; above that, nearly all massive halos host spheroidal galaxies [61, 62]. The lower limit in the $z$ integral should exceed the redshift of detectable birefringence sources. There is a trade-off between the galaxy number and the accumulation of birefringence along the line of sight. A more sophisticated calculation considering birefringence sources interspersed amongst the galaxies throughout all redshifts is model-dependent and is left for a future study.
The highest sensitivity achievable with data in the near future will likely come from correlating polarized SKA radio galaxies with LSST shape data (or using resolved SKA radio shapes). Similar to Ref. [63], we define Wide, Deep, and Ultradeep reference surveys, which have $\sigma_{I}=1\,\mathrm{\mu Jy}$, $0.2\,\mathrm{\mu Jy}$, and $0.05\,\mathrm{\mu Jy}$, and cover $1000\,\mathrm{deg}^{2}$, $20\,\mathrm{deg}^{2}$, and $1\,\mathrm{deg}^{2}$, respectively, and consider SKA-MID Band 5a (4.64.64.64.68.58.58.5\,8.5GHz [64]). Including galaxies whose integrated $4.8\,\mathrm{GHz}$ intensities have ${\rm SNR}>5$, we estimate the number of measurable spiral galaxies at $1.50.2<z<$ 1.5 to be $N=1.6\times 10^{6}$, $2.1\times 10^{5}$ and $2.8\times 10^{4}$ for each survey, respectively. For $1.50.8<z<$ 1.5, the numbers are $N=2.8\times 10^{5}$, $9.5\times 10^{4}$ and $1.7\times 10^{4}$, respectively (see App. D for calculation details).
For most detectable galaxies the polarization measurement error significantly contributes to the variance of the estimator $\hat{\alpha}$. Those galaxies collectively provide a non-negligible amount of information to reduce estimation noise. Consider $N$ galaxies in a small sky patch ${\rm d}\Omega$, each with a rotation estimator $\hat{\alpha}_{i}$. A linear combination $\hat{\alpha}_{\mathrm{best}}\equiv\sum^{N}_{i=1}w_{i}\,\hat{\alpha}_{i}$ with weights $w_{i}=\expectationvalue{\hat{\alpha}_{i}^{2}}^{-1}/\sum_{j}\expectationvalue{\hat{\alpha}_{j}^{2}}^{-1}$ gives an unbiased minimum-variance estimator. For different galaxies, $\expectationvalue{\hat{\alpha}_{i}^{2}}$ only depends on the unpolarized radio SNR, which in turn only depends on $z$ and $L_{4.8}$. Let us denote it as $\expectationvalue{\hat{\alpha}^{2}}(x)$, where $x=(z,\,\log L_{4.8})$.
With Eq. (17) rewritten as $N={\rm d}\Omega\,\int\Phi(x)\,{\rm d}x$, the weights are
@@
where $\Phi(x)$ quantifies the galaxy distribution with redshift and luminosity as defined in Eq. (43) in App. C. This leads to
@@
Accounting for the SNR distribution is therefore equivalent to having $N$ identically measured galaxies, each with an effective estimator variance
@@
While the $x$-integration formally depends on the choice of the threshold $S_{\nu}^{\mathrm{min}}$ as in Eq. (17), the integral approaches a minimal value as $S_{\nu}^{\mathrm{min}}$ decreases, until saturation at $S_{\nu}^{\mathrm{min}}/\sigma_{I}\lesssim 5$.
Fig. 3 shows the noise angular power spectrum of the rotation estimator, compared with sensitivity curves for current and projected CMB surveys [20]. SKA will enable a noise power spectrum comparable to or even lower than that of CMB-S4 for $L\gtrsim 5$, but probe a more limited $z$ range of birefringence sources. The noise spectra are compared to the maximal signal allowed by Planck [20] due to ultralight axion strings [13] approximated by the loop crossing model [65].
@@
Discussion — Using Eq. (9), rotation can be estimated at discrete galaxy locations $\{\hat{\alpha}_{i}\}_{i}$, which is suited for real-space approaches that exploit the non-Gaussian statistics of $\alpha(\bm{\mathbf{\hat{n}}})$. This is applicable to birefringence from axion strings [66, 67].
The reduced intrinsic noise ($5^{\circ}$$15^{\circ}$) of our carefully constructed estimator compared to the spread of the misalignment angle ($26^{\circ}$) is promising for reducing shot noise and bias due to intrinsic alignment in cosmic shear estimation [30].
Measuring birefringence in the CMB has the advantage of being sensitive to all sources up to $z\approx 1100$. Spiral galaxies do not effectively probe birefringence sources at $z\gtrsim 1.5$. Nonetheless, galaxies present an opportunity for extracting the $z$-dependence of the birefringence sources. One could compare CMB rotation estimation with that from radio galaxies, or compare galaxy estimation between different $z$ bins. For example, if axion strings exist for much of the cosmic history, this would enable the measurement of their size and number evolution with $z$ [65]. An axion field that is homogeneous but dynamic for $z\lesssim 1100$ [40] could also show up as $z$-dependent from galaxy-based tomographic estimation, complementing recently proposed tomography methods for the CMB [68, 69, 70, 71]. Furthermore, any detection of birefringence through the CMB [14] can be validated with galaxies. Foreground systematics , e.g. from the MW, would show up in every redshift bin and could thus be identified and subtracted.
Conclusion — We have developed an unbiased minimum-variance quadratic estimator to detect anisotropic cosmic birefringence from correlation between radio polarization and shape in spiral galaxies. This offers an independent technique that complements CMB probes. This will be enabled by a synergy between upcoming large radio galaxy surveys such as SKA and optical imaging surveys such as LSST. The noise floor of the polarization rotation angular power spectrum of the galaxy-based method using SKA will be better than that of CMB-S4. We plan to test the empirical correlations of [24] on cosmological galaxy samples using data from existing smaller radio surveys [36, 48, 43].
@@ Acknowledgements.
We thank Neal Dalal, Kangning Diao, Marc Kamionkowski, Dustin Lang, Christopher McKee, Jessie Muir, Kendrick Smith, Eiichiro Komatsu, Adrian Lee, Uroš Seljak, and Zack Li for useful discussions and comments. L.D. acknowledges research grant support from the Alfred P. Sloan Foundation (Award Number FG-2021-16495), from the Frank and Karen Dabby STEM Fund in the Society of Hellman Fellows, and from the Office of Science, Office of High Energy Physics of the U.S. Department of Energy (Award Number DE-SC-0025293). S.F. is supported by Lawrence Berkeley National Laboratory and the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade.
@@ Appendix A Quadratic estimator
Our goal is to estimate the polarization rotation field $\hat{\alpha}(\mathbf{n})$ using $X$ and $Y$. Since $X$ and $Y$ are computed from the ellipticity, one may be concerned about weak lensing effects. For CMB anisotropies, quadratic estimators for cosmic shear and for polarization rotation are orthogonal to each other to first order in each of the effects due to their opposite parity [20]. We expect this first-order non-contamination to hold for our birefringence estimator. Cosmic shear induces a white noise on $Y\approx p\,|\epsilon|\,\gamma=0.002\,(p/0.1)\,(\gamma/0.02)\,|\epsilon|$ where $\sqrt{\langle\gamma^{2}\rangle}\approx 0.02$ at $z\sim 1$ [72], which is smaller than the intrinsic $\sigma_{Y}=0.004$ (Eq. (12)). With both weak lensing and galaxy intrinsic shape alignment, a correlated noise contribution to $\hat{\alpha}$ arises, but with a magnitude significantly smaller than the noise floor derived here. We will therefore disregard weak lensing below.
We first comment on the measurement error in polarization. The important quantities for our purpose are $\sigma_{I}$, the measurement error in the total intensity, and $\sigma_{Q,U}$, the measurement error in the Stokes parameters $Q$ and $U$, assumed to be equal. For a monochromatic electromagnetic wave described by complex electric field amplitudes $E_{x},E_{y}$, the Stokes parameters are given by
$I$ $=|E_{x}|^{2}+|E_{y}|^{2},$ (23)
$Q$ $=|E_{x}|^{2}-|E_{y}|^{2},$ (24)
[…]
In the absence of any signal, we model components of $E_{x},E_{y}$ as independent Gaussian distributions with zero mean and equal variance. The variances of the Stokes parameters satisfy
$\sigma_{Q}^{2}=\sigma^{2}_{U}=\sigma^{2}_{V}=f\,\sigma^{2}_{I},$ (27)
with $f=1$.
This derivation applies to coherent monochromatic noise. While $E_{x},E_{y}$ are Gaussian, the Stokes parameters are non-Gaussian. In the practical case of integration over a frequency band, Eqs. (23)–(26) are modified to statistical averages over frequency and time. The distributions of the Stokes parameters $I,Q,U,V$ approach Gaussian by the central limit theorem. In other words, the measurement noise of the Stokes parameters in a realistic measurement of radio polarization may be assumed to be Gaussian.
In practice, the presence of a signal predominant in $I$ increases $\sigma^{2}_{I}$ more than $\sigma^{2}_{Q,U,V}$. The precise error in radio polarization measurements varies depending on the beam profile, frequency, duration of observation, number of pointings, and whether model fitting is involved in polarization detection, generally with $f\lesssim 1$. For measurements of spiral galaxies by VLA, this ratio is less than $1/2$ (e.g. 0.45 at $4.8\,\mathrm{GHz}$ and 0.36 at $1.4\,\mathrm{GHz}$ [73]). In forecasting sensitivities of SKA-MID configurations, we adopt $f=1$, leading to a conservative forecast. Note that Ref. [74] assumes $f=1/2$ for SKA-MID at $1.4\,\mathrm{GHz}$.
Next, we assume that the uncertainty in $p=P/I=\sqrt{Q^{2}+U^{2}}/I$ is dominated by uncertainty in the numerator rather than the denominator, as the unpolarized intensity typically has a much higher SNR than the polarized emission. For realistic deep optical imaging surveys with shape measurements, the radio polarization measurement error dominates over the error in ellipticity measurement, an assumption we adopt. Polarization and shape noise components are assumed to be statistically independent from each other and from the polarization and noise signals.
With tildes denoting observed quantities with measurement noise, for each galaxy with total flux $I$,
[…]
$=\expectationvalue{(\tilde{U}^{2}\,\epsilon_{1}^{2}+\tilde{Q}^{2}\,\epsilon_{2}^{2}-2\,\tilde{Q}\,\tilde{U}\,\epsilon_{1}\,\epsilon_{2})/I^{2}}$
$=\expectationvalue{Y^{2}}+(\sigma_{U}^{2}\expectationvalue{\epsilon_{1}^{2}}+\sigma_{Q}^{2}\expectationvalue{\epsilon_{2}^{2}})/I^{2}$
$=\expectationvalue{Y^{2}}+f\,(\sigma_{I}^{2}/I^{2})\expectationvalue{|\epsilon|^{2}}.$ (28)
After a similar calculation for $\expectationvalue{\tilde{X}^{2}}$, we obtain the errors in $X$ and $Y$ including measurement error:
$\tilde{\sigma}_{X,Y}^{2}=\sigma_{X,Y}^{2}+f\,\frac{\sigma_{\epsilon}^{2}}{\mathrm{SNR}^{2}},$ (29)
where the total flux SNR enters the expression, and $\sigma_{\epsilon}=\sqrt{\expectationvalue{|\epsilon|^{2}}}\approx 0.3$ as in the main text.
@@
where the coefficients are to be determined.
Under a polarization rotation $\alpha$ at the galaxy’s location, the observed quantities are transformed:
$\tilde{X}\to\tilde{X}-2\alpha\,Y,\quad\tilde{Y}\to\tilde{Y}+2\alpha\,X,$ (31)
where the small angle approximation ($\alpha\ll 1$) is used, since large birefringence angles are observationally ruled out. Note that the transformation is proportional to quantities without measurement noises, since cosmic birefringence is applied before detection. By requiring that the estimator $\hat{\alpha}$ be unbiased, namely $\expectationvalue{\hat{\alpha}}=\alpha$, we have
@@
Using these, $u$ can be expressed in terms of $u^{\prime}$ and $v^{\prime}$, and $v$ can be expressed in terms of $w$. It remains to minimize the variance of the estimator, $\expectationvalue{\hat{\alpha}^{2}}$, under the null hypothesis, with respect to $u^{\prime}$, $v^{\prime}$, and $w$.
To enable analytic calculations, we approximate $X$ as a Gaussian random variable with mean $\overline{X}$ and variance $\sigma_{X}^{2}$, and $Y$ a Gaussian random variable with zero mean and variance $\sigma_{Y}^{2}$. Since $X$ and $Y$ are only approximately Gaussian, this gives a nearly optimal quadratic estimator. The true optimal estimator could be obtained by numerically minimizing $\expectationvalue{\hat{\alpha}^{2}}$ given empirical distributions for $Ip$, $|\epsilon|$, and $\Delta\theta$.
After some algebra, we obtain the claimed result in Eq. (9):
@@ Appendix B Maximum likelihood approach
where $\mathbf{N}_{i}$ is the measurement error of $\mathbf{X}_{i}$; $\overline{\mathbf{X}}_{i}$ is the mean of $\mathbf{X}$ over an ensemble of galaxies; $C_{\mathbf{X}_{i}}$ is the covariance matrix of the intrinsic, unrotated quantities $\mathbf{X}_{i}$ (which in general may depend on the type or other properties of the $i$-th galaxy); $C_{\mathbf{N}_{i}}$ is the covariance matrix of the measurement noise given by the second term in Eq. (11), meaning $C_{\mathbf{N}_{i}}=(\sigma_{\epsilon}^{2}/2)\,\mathrm{SNR}_{i}^{2}\operatorname{diag}(1,1)$; and $M(\alpha(\bm{\mathbf{\hat{n}}}_{i}))$ is the effect of polarization rotation on the intrinsic quantities. Eq. (35) assumes that each $\mathbf{X}_{i}$ is a random Gaussian variable, the same approximation we used in the main text, meaning $C_{\mathbf{X}_{i}}=\operatorname{diag}(\sigma_{X_{i}}^{2},\sigma_{Y_{i}}^{2})$. Our goal is then to obtain an estimator $\hat{\alpha}_{\mathrm{ML}}(\bm{\mathbf{\hat{n}}})$ which maximizes $\ln\mathcal{L}$ given a set of observations $\{\mathbf{X}_{i}\}$.
@@
$\ln\mathcal{L}[\alpha;\mathbf{X},\mathbf{N}=\mathbf{N}_{\mathrm{max}}]=-\frac{1}{2}(\mathbf{X}-\mathbf{M}^{-1}\overline{\mathbf{X}})^{T}(C_{\mathbf{N}}+\mathbf{M}^{T}C_{\mathbf{X}}\mathbf{M})^{-1}(\mathbf{X}-\mathbf{M}^{-1}\overline{\mathbf{X}}).$ (38)
@@
We insert this expansion into Eq. (38) and take the derivative with respect to $\alpha$, keeping terms up to linear order in $\alpha$:
$\frac{\partial\mathcal{L}}{\partial\alpha}=\frac{2}{\tilde{\sigma}_{X}^{2}\,\tilde{\sigma}_{Y}^{2}}\left[\tilde{\sigma}_{Y}^{2}\,\overline{X}\,\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}\right]-\frac{4\,\alpha}{\tilde{\sigma}_{X}^{2}\,\tilde{\sigma}_{Y}^{2}}\left[\tilde{\sigma}_{Y}^{2}\,\overline{X}\tilde{X}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\left(\tilde{X}^{2}-\tilde{Y}^{2}\right)\right].$ (40)
Setting this to zero results in the maximum-likelihood estimator
$\hat{\alpha}_{\mathrm{ML}}=\frac{\tilde{\sigma}_{Y}^{2}\,\overline{X}\,\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}}{2\left[\tilde{\sigma}_{Y}^{2}\,\overline{X}\tilde{X}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\left(\tilde{X}^{2}-\tilde{Y}^{2}\right)\right]}.$ (41)
@@
$\hat{\alpha}_{\mathrm{ML}}=\frac{\tilde{\sigma}_{Y}^{2}\,\overline{X}\,\tilde{Y}+(\sigma_{X}^{2}-\sigma_{Y}^{2})\,\tilde{X}\,\tilde{Y}}{2\left[\tilde{\sigma}_{X}^{2}\,\overline{X}^{2}+(\sigma_{X}^{2}-\sigma_{Y}^{2})^{2}\right]}.$ (42)
The quadratic estimator in Eq. (9) is thus recovered. When $\alpha=0$, the variance $\expectationvalue{\hat{\alpha}_{\mathrm{ML}}^{2}}$ is therefore also given by Eq. (10).